OpenStudy (anonymous):

Can you please help me here? Trigo behaviors Q: State the period, amplitude, max/min values, range, domain, horizontal phase shift and vertical displacement.

OpenStudy (experimentx):

I Guess \( y = A\cos(wt + \phi) + C\) A will be your amplitude so, A=3 \(\omega \) will be your (angular *in physics we usually call that) frequency Period = \(\omega/2\pi\) \(\phi\) will be initial phase C will be your vertical shift

OpenStudy (experimentx):

*period will be in pi

OpenStudy (experimentx):

looks like i missed something.

OpenStudy (shayaan_mustafa):

may I help here?

OpenStudy (anonymous):

I would love that:)

OpenStudy (experimentx):

Not sure ... it could be sine ... since sin is also harmonic. period will be 1pi

OpenStudy (shayaan_mustafa):

experimentX is right. and your too open2study

OpenStudy (experimentx):

@Shayaan_Mustafa continue ..

OpenStudy (shayaan_mustafa):

ok thnx..

OpenStudy (shayaan_mustafa):

1)Write your equation as, \[\Large y=-3\cos[2(x-(\frac{30}{2})]\] look how i wrote the original equation. \[A=-3\]\[Period=\frac{2 \pi}{B}=\frac{2 \pi}{2}=\pi\]\[\text{phase shift}=\frac{C}{B}=\frac{\pi}{12}=15^{o}\]\[\text{vertical shift}=+1\]

OpenStudy (shayaan_mustafa):

general form